# Federated Nonlinear System Identification

### v2

@@ Abstract
1 Introduction
Dynamical system models describe how a system evolves over time based on its current state, control inputs, and external disturbances. These models are central to fields including control theory, physics, and robotics (Willems, 1989). A common class of such systems is linear time-invariant (LTI) systems, where the next state and output are linear functions of the current state and input, governed by parameters that remain constant over time (Sibai & Mitra, 2018). In contrast, nonlinear systems allow these dependencies to be governed by nonlinear functions, offering more accurate representations for many real-world systems with complex or nonstationary behavior (Guckenheimer & Holmes, 2013; Jiang & Lai, 2019).
@@ 1 Introduction
System identification is the process of learning dynamical models from input and output data. In this way, one can learn the dynamics of a system in a data-driven way, without explicit physical derivations. It is often used to learn dynamical systems in the real world, where exact dynamics are unknown, with applications including robotics, fluid dynamics, and physics (Sarkar et al., 2021; Venkatesh & Dahleh, 2002). Traditional approaches for system identification typically assume access to centralized datasets. However, in modern deployments—such as multi-robot systems, sensor networks, and connected vehicles—data is naturally distributed across many devices. Transmitting all data to a central location is often impractical due to privacy, bandwidth, or energy limitations.
To address these challenges, we consider federated system identification, which is emerging as a privacy-preserving alternative. In this paradigm, multiple clients collaboratively learn a model without sharing raw data, communicating only model updates. Prior work has explored federated identification for LTI systems (Wang et al., 2023), demonstrating convergence benefits that scale with the number of participating clients. However, this line of work is restricted to linear models, which fundamentally limits its applicability.
@@
In this paper, we study the problem of federated nonlinear system identification. We consider a setting where $M$ distributed clients each observe trajectory data from potentially different nonlinear dynamical systems that nonetheless belong to the same underlying family (see Figure 1). Dynamical systems at clients exhibit heterogeneity, modeled via a heterogeneity parameter $\epsilon$. This setup captures realistic scenarios such as learning dynamics across a fleet of heterogeneous robots or drones—each operating under different conditions due to hardware differences, controller tunings, or external disturbances. The goal is to identify a shared nonlinear model that generalizes across these clients while respecting local variability and preserving data privacy.
1.1 Key Contributions
Figure 2: Impact of client count and heterogeneity on federated nonlinear system identification. (a) Estimation error versus number of clients $M$ at fixed trajectories per client $N_{i}=10$ and heterogeneity bound $\epsilon=0.01$; (b) Estimation error versus $M$ for the same $N_{i}$ with $\epsilon=5$, illustrating increased model-mismatch tolerance; (c) Estimation error as a function of $\epsilon$ for fixed $M$ and $N_{i}$, showing sensitivity of global model accuracy to client-level parameter variations.
@@ 1.1 Key Contributions
  • Nonlinear framework. We consider linearly parameterized nonlinear dynamical systems, with a focus on the widely used piecewise affine (PWA) model for representing nonlinear dynamics (Mania et al., 2022; Musavi et al., 2024). In PWA models, state transitions are expressed as linear functions of nonlinear feature embeddings of state-input pairs. While prior work has explored federated system identification of linear dynamical systems, to the best of our knowledge, this work develops the first federated learning framework for system identification of PWA-based nonlinear dynamical systems.
  • Theoretical analysis. We theoretically analyze the convergence of the federated nonlinear system identification problem. Our results show that clients benefit significantly by collaborating through federated learning. Particularly, we show that the convergence error decreases as $\mathcal{O}(1/\sqrt{M})$, where $M$ is the number of clients, thus having significant improvement in convergence as more clients collaborate. Additionally, our results theoretically characterize how heterogeneity across clients impacts the convergence error.
@@
  • Experimental analysis. To corroborate our theoretical results, we conduct extensive experiments analyzing the convergence behavior of our approach as a function of the total number of participating clients, the number of local samples at each client, as well the heterogeneity of dynamical systems across clients. We preview some results in Figure 2, which are based on real-world pendulum dynamics. Complete details are provided in Section 4.
@@
  • Algorithm validation. Our experiments demonstrate that our federated approach achieves improved convergence rate compared to a single-client system, while offering the additional advantage of client data privacy.
2 Related Works
2.1 System Identification
@@
Ljung (1998) provides an overview of mathematical methods for building models of dynamical systems using input and output signal measurements, focusing on how parameter estimation techniques can be applied to characterize system behavior in both time and frequency domains. This approach is instrumental in transforming observed data into predictive models, supporting applications in signal processing and control. Jedra & Proutiere (2020) present a robust finite-time error bound for identifying stable linear systems using the ordinary least-squares (OLS) estimator. Simchowitz et al. (2018) also analyze OLS identification from a single trajectory for stable and marginally stable linear time-invariant systems, using the block-martingale small-ball method to address mixing dependencies and establish nearly optimal estimator rates for a wide class of linear time series. While deriving tight, regime-specific finite-time error rates for OLS identification of general linear systems, Sarkar & Rakhlin (2019) establish that OLS is optimally consistent only under specific regularity conditions, unifying previous results across all spectral regimes. However, this approach struggles with nonlinear or complex signal structures, often requiring significant computational resources and large datasets to achieve reliable identification.
2.2 Federated Learning
@@
Federated learning is a machine learning paradigm which allows clients to collaboratively train a model without sharing raw data (McMahan et al., 2017). In principle, each client possesses a local dataset and sends weight updates to the server. The server then aggregates these weights together to update the global model. This approach is very useful to train models on sensitive or private data. Truex et al. (2019) propose a hybrid federated learning approach that combines differential privacy and secure multiparty computation to enhance privacy protection while maintaining high model accuracy in collaborative machine learning settings. The paper presents communication-efficient federated learning strategies through structured low-dimensional updates and compressed sketched updates, achieving a substantial reduction in communication overhead while preserving model accuracy (Konečnỳ et al., 2016). Li et al. (2020) introduce the FedProx framework, which adds a proximal term to FedAvg and provides convergence guarantees to achieve more robust and stable training in the presence of statistical and system heterogeneity across clients. Federated learning architectures can use a peer-to-peer approach (Yang et al., 2019). However, our discussion centers on the traditional server-client model, in which a central server periodically aggregates updates from multiple clients. While most federated learning works are concerned with learning deep neural networks, our work focuses on understanding the theory and implementation of federated system identification of nonlinear dynamical systems.
2.3 Federated System Identification
@@
The goal of the federated system identification problem is to determine a shared estimate of the system matrices, denoted as $\bar{\theta}$, that yields a small estimation error relative to each client $i$’s true system parameters $\theta^{(i)*}$. Clients $i$ do not share their private data with the server, and each client independently estimates its unknown system dynamics. Wang et al. (2023) formulated an FL framework for system identification of linear system models across heterogeneous clients with improved sample efficiency. They show that more clients improve the model convergence rate and propose the FedSysID algorithm. Their work was restricted to synthetic linear systems, which limits applicability to real-world nonlinear dynamical systems. This limitation motivates our proposal of federated nonlinear system identification, as detailed in the following sections.
3 Federated Nonlinear System Identification
@@
3.1 Preliminaries
The formal objective of the federated system identification problem is to estimate a model that accurately captures the collective behavior of the participating systems. Specifically, suppose that for each client $i\in[M]$, we have the following linearly parameterized nonlinear dynamics:
@@ 3.1 Preliminaries
where $x_{t}^{(i)}$, $u_{t}^{(i)}$, and $w_{t}^{(i)}$ denote the state, control, and disturbance respectively. Furthermore, $\phi:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{n}}\to\mathbb{R}^{n_{\phi}}$ denotes a vector of known nonlinear mappings (Mania et al., 2022; Musavi et al., 2024).
In the following, we recall two key definitions from Musavi et al. (2024) that will be used in our assumptions.
Definition 1.
(Semi-continuous distribution). A probability distribution $\mathbb{P}$ is semi-continuous if there does not exist a set $E$ with Lebesgue measure zero such that $\mathbb{P}(E)=1$.
Definition 2.
@@
We next provide four assumptions for the local dynamical systems at clients (Musavi et al., 2024) (we omit client indexing for brevity):
@@
  • Assumption 2: Noise $w_{t}$ is i.i.d., and follows a semi-continuous, mean-zero distribution with positive semi-definite covariance matrix $\sum_{w}\succeq\sigma_{w}^{2}I_{n_{x}}$.
  • Assumption 3: Input $u_{t}$ is i.i.d following a semi-continuous, mean-zero distribution with positive semi-definite covariance matrix $\sum_{u}\succeq\sigma_{u}^{2}I_{n_{x}}$.
@@
Additionally, inspired by Wang et al. (2023), we consider the following assumption of bounded system heterogeneity across clients.
@@
3.2 Framework Convergence Rate
A least-squares estimate is meaningful only if the design matrix built from each client’s regressor vectors is well conditioned. In linear system identification this role is played by a persistent-excitation assumption. For linearly parameterized nonlinear dynamics, we recover a similar property through the block-martingale small-ball (BMSB) condition. Lemma 1 is used to show that the analytic feature map and new inputs guarantee the required excitation.
Lemma 1 (BMSB for open-loop systems).
Let each client $i\in[M]$ run open-loop inputs $u^{(i)}_{t}=\eta^{(i)}_{t}$. Let client $i$ collect a trajectory of length $T$. For every $t=0,1,\dots,T-1$, define the filtration
@@ 3.2 Framework Convergence Rate
Under Assumptions 1–4 there exist constants $s_{\phi}>0$ and $p_{\phi}\in(0,1)$ (as defined by Musavi et al. (2024)) such that, for every client $i$ and every unit vector $v\in\mathbb{S}^{n_{\phi}-1}$,
@@
Less formally, each new regressor $\phi(x^{(i)}_{t},u^{(i)}_{t})$ has (conditioned on the past) at least probability $p_{\phi}$ of having length $s_{\phi}$ in every direction, hence the excitation we desired.
@@
Theorem 1 (Finite-sample error).
@@
Remark 1.
The first term in the convergence error decreases as $1/\sqrt{T\sum_{i=1}^{M}N_{i}}$, where $\sum_{i=1}^{M}N_{i}$ is the total number of trajectories across all clients. Furthermore, since the number of trajectories $N_{i}$ at client $i$ is a constant for a given trajectory length $T$, the error decreases as $\mathcal{O}(1/\sqrt{M})$.
Remark 2.
For smaller $\epsilon$, the underlying system dynamics at the clients are more similar, enabling FL clients to effectively use data from all participants. This enhances the performance of each client, particularly achieving a convergence speedup of $\sqrt{M}$.
3.3 FNSysId
The core workflow of FNSysId implementation is formalized in Algorithm 1, which shows how clients perform local computations and synchronize model updates with the central server during each communication round.
Algorithm 1 FNSysId
Initialize the server with $\bar{\theta}_{0}$, and learning rate $\alpha$ ;
Initialize each client $i\in[M]$ with $\theta_{0,0}^{{(i)}}=\bar{\theta}_{0}$;
For each global iteration $r=0,1,\ldots,R-1$ do
@@ 3.3 FNSysId
  For each client $i\in[M]$ in parallel do
   $\theta_{r+1}^{{(i)}}=\texttt{ClientUpdate}(i,\bar{\theta}_{r},K_{i})$
   send $\theta_{r+1}^{{(i)}}$ back to the server
  end for
@@
  update $\bar{\theta}_{r+1}=\frac{1}{M}\sum_{i=1}^{M}\theta_{r+1}^{{(i)}}$ and send $\bar{\theta}_{r+1}$ to each client
end for
Return $\bar{\theta}_{R}$
In our framework, we assume all clients $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{M}\}$ actively participate in every global iteration $r\in[R]$. Each client performs $K_{i}$ local updates independently using its own local data before any communication with the central server (line 6), where ClientUpdate iteratively executes:
$\theta_{r,k}^{{(i)}}=\;$ $\theta_{r,k-1}^{{(i)}}+\alpha\left({y}^{{(i)}}-\theta_{r,k-1}^{{(i)}}{X}^{{(i)}}\right)({X}^{{(i)}})^{\top},\quad k=1,2,\ldots,K_{i}.$ (7)
This approach reduces communication overhead by allowing multiple updates locally. After completing these updates, the clients send their updated local models $\bar{\theta}_{r+1}^{(i)}$ to the server. The server then averages all updates to obtain a new global model $\bar{\theta}_{r+1}$, which is subsequently broadcast back to all clients to synchronize (line 10). This efficiently balances local computation and communication, addressing potentially limited communication capabilities between clients.
4 Experiments
We conducted experiments on a range of systems, encompassing both synthetic benchmarks and real-world dynamical systems. In the following, we provide details regarding datasets employed, hyperparameter configurations, and the corresponding results.
4.1 Data
Conducting experiments on both synthetic and real-world system data is essential to comprehensively evaluate the proposed methodology. Synthetic datasets offer controlled environments enabling precise assessment of model accuracy and convergence behavior. In contrast, real-world datasets provide insights into the method’s practical applicability, robustness to noise, and generalization capabilities under complex, uncontrolled conditions. This dual evaluation ensures both theoretical soundness and empirical relevance.
4.1.1 Synthetic Data
To evaluate the performance and efficiency of FNSysId, we conduct numerical experiments involving $M$ clients, each associated with a discrete-time dynamical system characterized by $n=3$ states and $p=2$ inputs. The initial state, input signal, and process noise are all sampled from zero-mean distributions with unit standard deviations. The trajectories were generated by augmenting the trajectory formulation presented by Wang et al. (2023) with a nonlinear term, as presented in the following equation.

### v3

# arXiv:2508.15025v3[cs.LG] 27 Feb 2026

@@ Abstract
I Introduction
Dynamical system models describe how a system evolves over time based on its current state, control inputs, and external disturbances. These models are important to fields such as control theory, physics, and robotics [1]. A common category of systems is linear time-invariant (LTI) systems, where both the next state and output are linear functions of the current state and input. These systems are governed by parameters that remain constant over time [2]. In contrast, nonlinear functions determine these relationships in nonlinear systems [3, 4].
@@ I Introduction
System identification is a data-driven approach to learn dynamical models from input and output data. It is often used to learn dynamical systems for which exact dynamics are unknown, with applications including robotics, fluid dynamics, physics and healthcare [5, 6, 7]. Traditional approaches for system identification assume centralized data. However, in modern applications, data is often distributed across many devices, and transmitting all data to a central server is often impractical due to privacy, bandwidth, or energy limitations. To address these challenges, we consider federated system identification, which is emerging as a privacy-preserving alternative. In this paradigm, multiple clients collaboratively learn a model without sharing raw data, communicating only model updates. Prior work has explored federated identification for LTI systems [8].
@@
In this paper, we study the problem of federated nonlinear system identification. We consider a setting where $M$ distributed clients each observe trajectory data from different nonlinear dynamical systems that belong to the same underlying family (see Figure 1). The dynamical systems we consider exhibit heterogeneity, modeled by the heterogeneity parameter $\epsilon$. Federated learning should be used for system identification when multiple similar systems share common structure and data cannot be centralized, enabling privacy-preserving collaborative learning. Results show that aggregating across more participating clients improves convergence and reduces individual client noise.
Our main contributions are as follows:
@@
  • Nonlinear framework. We consider linearly-parameterized nonlinear dynamical systems, with a focus on the widely used piecewise affine (PWA) model for representing nonlinear dynamics [9, 10]. In PWA models, state transitions are expressed as linear functions of nonlinear feature embeddings of state-input pairs. While prior work has explored federated system identification of linear dynamical systems, to the best of our knowledge, this work develops the first federated learning framework for system identification of PWA-based nonlinear dynamical systems.
[…]
@@
  • Experimental analysis. We conduct experiments analyzing the convergence behavior of our approach as a function of the total number of participating clients, the number of local samples at each client, as well the heterogeneity of dynamical systems across clients. Complete details are provided in Section IV.
@@
  • Algorithm validation. Our experiments demonstrate that our federated approach achieves an improved convergence rate compared to a single-client system.
II Related Works
II-A System Identification
@@
The paper [11] provides an overview of mathematical methods for building models of dynamical systems using input and output signal measurements, focusing on how parameter estimation techniques can be applied to characterize system behavior in time and frequency domains. This approach is instrumental in transforming observed data into predictive models, supporting applications in signal processing and control. System identification can be extended to linearly-parameterized nonlinear systems [10]. The authors in [12] present a robust finite-time error bound for identifying stable linear systems using the ordinary least-squares (OLS) estimator. The study in [13] also analyzes OLS identification from a single trajectory for stable and marginally stable linear time-invariant systems, using the block-martingale small-ball method to address mixing dependencies and establishes nearly optimal estimator rates for a wide class of linear time series. The work further provides tight, regime-specific finite-time error bounds for OLS identification of general linear dynamical systems. The framework in [14] establishes that OLS is optimally consistent only under specific regularity conditions, unifying previous results across all spectral regimes. This approach struggles with nonlinear or complex signal structures, often requiring significant computational resources and large datasets to achieve reliable identification.
II-B Federated Learning
@@
Federated learning is a machine learning paradigm which allows clients to collaboratively train a model while preserving data privacy [15]. Each client possesses a local dataset and sends weight updates to the server. The server then aggregates these weights and updates the global model. The method presented by [16] is a hybrid federated learning approach that combines differential privacy and secure multi-party computation to enhance privacy protection while maintaining high model accuracy. Communication-efficient federated learning strategies through structured low-dimensional updates and compressed sketched updates may achieve a substantial reduction in communication overhead while preserving model accuracy [17]. The work of[18] introduces the FedProx framework, which adds a proximal term to FedAvg and provides convergence guarantees to achieve more robust and stable training in the presence of statistical and system heterogeneity across clients. Federated learning architectures can use a peer-to-peer approach [19], however, our discussion focuses on the traditional server-client model, in which a central server periodically aggregates updates from multiple clients. Although most federated learning works are concerned with learning deep neural networks, our work focuses on understanding the theory and implementation of federated system identification of nonlinear dynamical systems.
II-C Federated System Identification
@@
The goal of federated system identification is to learn a shared estimate of the system parameters that yields a small estimation error relative to the true parameters of each client $i$’s true system parameters $\theta^{(i)*}$. Clients $i$ do not share their private data with the server, and each client independently estimates its unknown system dynamics. The paper [8] formulated an FL framework for system identification of linear system models across heterogeneous clients with improved sample efficiency, proposing the FedSysID algorithm and showing that more clients improve the model convergence rate. However, [8] was restricted to synthetic linear systems. The authors of [20] introduce Koopman learning as a method for dealing with nonlinear systems that have partially observable data. The setting is where multiple clients observe an identical nonlinear system, but the system lacks any active control ($u_{t})$, which is contrary to what we will consider in (1). These limitations motivate our work on federated nonlinear system identification.
III Federated Nonlinear System Identification
@@
III-A Preliminaries
The formal objective of the federated system identification problem is to estimate a model that accurately captures the collective behavior of the participating systems. Formally, first suppose that for each client $i\in[M]$, we have the following linearly-parameterized nonlinear dynamics:
@@ III-A Preliminaries
where $x_{t}^{(i)}$, $u_{t}^{(i)}$, and $w_{t}^{(i)}$ denote the state, control, and disturbance respectively. Furthermore, $\phi:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{\phi}}$ denotes a vector of known nonlinear mappings [9, 10].
In the following, we recall two key definitions from [10] that will be used in our assumptions. Refer to [10] for further clarification on these definitions.
Definition 1
(Semi-continuous distribution). A Euclidean probability distribution $\mathbb{P}$ is semi-continuous if there does not exist a set $E$ with Lebesgue measure zero such that $\mathbb{P}(E)=1$.
Definition 2
@@
We next provide four assumptions for the local dynamical systems at clients [10] (we omit client indexing for brevity).
@@
  • Assumption 2: Noise $w_{t}$ is i.i.d., and follows a semi-continuous, zero-mean distribution with positive semi-definite covariance matrix $\sum_{w}\succeq\sigma_{w}^{2}I_{n_{x}}$.
  • Assumption 3: Input $u_{t}$ is i.i.d following a semi-continuous, zero-mean distribution with positive semi-definite covariance matrix $\sum_{u}\succeq\sigma_{u}^{2}I_{n_{x}}$.
@@
Additionally, inspired by [8], we consider the following assumption of bounded system heterogeneity across clients.
@@
III-B Convergence Rate
A least-squares estimate is meaningful only if the design matrix built from each client’s regressor vectors is well conditioned. In linear system identification this role is played by a persistent‐excitation assumption. For linearly-parameterized nonlinear dynamics, we recover a similar property through the block‑martingale small‑ball (BMSB) condition (Definition 4). Lemma 1 is used to show that the analytic feature map and new inputs guarantee the required excitation.
Lemma 1 (BMSB for open‑loop systems)
Let each client $i\in[M]$ run open‑loop inputs $u^{(i)}_{t}=\eta^{(i)}_{t}$. Let client $i$ collect a trajectory of length $T$. For every $t=0,1,\dots,T-1$, define the filtration
@@ III-B Convergence Rate
Under Assumptions 1–4 there exist constants $s_{\phi}>0$ and $p_{\phi}\in(0,1)$ (as defined by [10]) such that, for every client $i$ and every unit vector $v\in\mathbb{S}^{n_{\phi}-1}$,
@@
Each new regressor $\phi(x^{(i)}_{t},u^{(i)}_{t})$ has (conditioned on the past) at least probability $p_{\phi}$ of having length $s_{\phi}$ in every direction, hence the excitation we desired.
@@
Theorem 1 (Finite‑sample error)
@@
Remark 1
The first term in the convergence error decreases as $1/\sqrt{T\sum_{i=1}^{M}N_{i}}$, where $\sum_{i=1}^{M}N_{i}$ is the total number of trajectories across all clients. Furthermore, since the number of trajectories $N_{i}$ at client $i$ is a constant for a given trajectory length $T$, the error decreases as $\tilde{\mathcal{O}}(1/\sqrt{M})$.
Remark 2
For smaller $\epsilon$, the underlying system dynamics at the clients are more similar, enabling FL clients to effectively use data from all participants. This enhances the performance of each client, particularly achieving a convergence speedup of ${\approx}\sqrt{M}$.
III-C FNSysId
The FNSysId  [^2] The source code is available for download - https://github.com/OMKARTT/FNSysId implementation is formalized in Algorithm 1, which shows how clients perform computations and synchronize model updates with the central server during each communication round. This is the standard FL algorithm.
In our framework, we assume all clients $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{M}\}$ actively participate in every global iteration $r\in[R]$.
Each client performs $K_{i}$ local updates independently using its own local data before any communication with the central server (line 6), where NonlinearClientUpdate iteratively executes:
$\theta_{r,k}^{{(i)}}=\;$ $\theta_{r,k-1}^{{(i)}}+\alpha\left({y}^{{(i)}}-\theta_{r,k-1}^{{(i)}}{X}^{{(i)}}\right)(X^{(i)})^{T},$ (7)
where $X^{(i)}$ is the input data and $y^{(i)}$ is the true system. $k=1,2,\ldots,K_{i}.$ Note that this is a restatement of FedAvg.
This approach reduces communication overhead by allowing multiple updates locally. After completing these updates, the clients send their updated local models $\bar{\theta}_{r+1}^{(i)}$ to the server. The server then averages all updates to obtain a new global model $\bar{\theta}_{r+1}$, which is subsequently broadcast back to all clients to synchronize (line 11). This efficiently balances local computation and communication, addressing potentially limited communication capabilities between clients.
Algorithm 1 FNSysId: Federated Nonlinear System Identification
1:Initialize server with global parameter $\bar{\theta}_{0}$ and learning rate $\alpha$ ;
2:Initialize each client $i\in[M]$ with local parameter $\theta^{(i)}_{0,0}=\bar{\theta}_{0}$ ;
3:for global iteration $r=0,1,\ldots,R-1$ do
@@ III-C FNSysId
5:  for each client $i\in[M]$ in parallel do
6:    $\theta^{(i)}_{r+1}=\texttt{NonlinearClientUpdate}\big(i,\bar{\theta}_{r},K_{i}\big)$ (7)
7:    Send $\theta^{(i)}_{r+1}$ to server
8:  end for
@@
10:  $\bar{\theta}_{r+1}=\frac{1}{M}\sum_{i=1}^{M}\theta^{(i)}_{r+1}$
11:  Broadcast $\bar{\theta}_{r+1}$ to all clients
12:end for
13:Return global estimate $\bar{\theta}_{R}$
Corollary 1
Using the same setup as Theorem 1 and for all communication rounds $R\geq 1$, the output $\bar{\theta}_{R}$ given by the FNSysId algorithm with the FedAvg update rule satisfies:
$\max_{i\in[M]}\|\bar{\theta}_{R}-\theta^{(i)*}\|_{2}\leq\mathcal{O}\left(\frac{1}{KR}+\frac{1}{\sqrt{\sum_{i=1}^{M}N_{i}T}}+\epsilon\right).$
Proof:
Because the system studied in this paper is linearly-parameterized, the proof directly follows from [8, Corollary 4]. ∎
IV Experiments
IV-A Data
In this section we describe the datasets used in our experiments.
IV-A1 Synthetic Data
To evaluate the performance and efficiency of FNSysId, we conduct experiments involving $M$ clients, each associated with a discrete-time dynamical system characterized by $n=3$ states and $p=2$ inputs. The initial state, input signal, and process noise are all sampled from zero-mean distributions with unit standard deviations. The trajectories were generated by augmenting the trajectory formulation presented by [8] with a nonlinear term, as presented in the following equation.